OpenStudy (anonymous):

Which of the following is the simplified form of x+7/2x+12 + 6/x^2-36?

OpenStudy (lgbasallote):

x+7/2(x+6) + 6/(x+6)(x-6) x+7(x-6) + 6 all over (X+6)(x-6) (x^2 + 13x -36)/(x+6)(x-6)

OpenStudy (anonymous):

not a choice

OpenStudy (lgbasallote):

it isn't???

OpenStudy (anonymous):

no sorry

OpenStudy (lgbasallote):

oh yeah...forgout about the 2 there :)))

OpenStudy (lgbasallote):

x+7/2(x+6) + 6/(x+6)(x-6) x+7(x-6) + 12 all over 2(X+6)(x-6) (x^2 + 13x -30)/2(x+6)(x-6)

OpenStudy (lgbasallote):

that in the choices?

OpenStudy (mertsj):

\[\frac{x+7}{2x+12}\times\frac{6}{x^2-36}\]

OpenStudy (anonymous):

no its not

OpenStudy (mertsj):

\[\frac{x+7}{2(x+6)}\times\frac{6}{(x-6)(x+6)}\]

OpenStudy (anonymous):

x+5/2(x-6) or x-5/2(x+6) or x-5/2)x-6) or x+5/2(x+6)

OpenStudy (mertsj):

\[\frac{(x+7)(x-6)}{2(x+6)(x-6)}+\frac{12}{2(x-6)(x+6)}\]

OpenStudy (mertsj):

\[\frac{x^2+x-42+12}{2(x-6)(x+6)}\]

OpenStudy (mertsj):

\[\frac{x^2+x-30}{2(x-6)(x+6)}\]

OpenStudy (mertsj):

\[\frac{(x+6)(x-5)}{2(x-6)(x+6)}\]

OpenStudy (lgbasallote):

oh...i see...my answer wasn't simplified...x -5/2(x-6)...that right now?

OpenStudy (mertsj):

\[\frac{x-5}{2(x-6)}\]